# On the Learnability of Watermarks for Language Models

### v2

@@ Abstract
Watermarking of language model outputs enables statistical detection of model-generated text, which has many applications in the responsible deployment of language models. Existing watermarking strategies operate by altering the decoder of an existing language model, and the ability for a language model to directly learn to generate the watermark would have significant implications for the real-world deployment of watermarks. First, learned watermarks could be used to build open models that naturally generate watermarked text, allowing for open models to benefit from watermarking. Second, if watermarking is used to determine the provenance of generated text, an adversary can hurt the reputation of a victim model by spoofing its watermark and generating damaging watermarked text. To investigate the learnability of watermarks, we propose watermark distillation, which trains a student model to behave like a teacher model that uses decoding-based watermarking. We test our approach on three distinct decoding-based watermarking strategies and various hyperparameter settings, finding that models can learn to generate watermarked text with high detectability. We also find limitations to learnability, including the loss of watermarking capabilities under fine-tuning on normal text and high sample complexity when learning low-distortion watermarks. [^1] We publicly release code and trained model weights at https://github.com/chenchenygu/watermark-learnability.
@@ 1 Introduction
As large language models (LLMs) become more capable and widely used, watermarking language model outputs becomes increasingly important to mitigate potential harms and misuses of LLMs. Watermarking enables statistical detection of model-generated text, which enables enforcing policies around LLM usage, e.g., removing LLM-generated disinformation from social media platforms or detecting academic dishonesty. Another proposed use case of watermarking is identifying text provenance, i.e., tracing a text back to the specific LLM that generated it (Abdelnabi & Fritz, 2021; Kuditipudi et al., 2023).
@@
Recent works have demonstrated that it is possible for an LLM provider to inject specific, known watermark patterns into text using specialized decoding algorithms (Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023), but little is known about whether these watermarks are learnable by a model. The learnability of watermarks has significant implications for the real-world deployment of watermarks, as it could enable downstream applications and adversarial attacks.
@@
In this work, we study the learnability of watermarks by studying weights-based watermarking, which involves learning parameters for a language model that cause it to generate watermarked text under its natural sampling distribution, without using a special decoding-time watermark algorithm. Our investigation is motivated by its relevant implications for two applications: developing watermarking for open language models and the possibility of watermark spoofing attacks.
Existing watermarking methods depend upon using a specialized decoding algorithm. This makes them too brittle for open models, as users will use a variety of decoding algorithms due to user choice, different inference infrastructures and environments, etc. Weights-based watermarking removes the reliance on a specialized decoder and is robust to changes in decoding algorithm, making it a necessary step towards building open models that generate watermarked text using only their weights.
Watermark spoofing attacks are hypothetical attacks proposed in Sadasivan et al. (2023) where an adversary outputs text that contains the watermark signal from a victim model. If watermarking is used to identify text provenance, then such an attack can attribute harmful behavior to and damage the reputation of the victim model. The ability to learn watermarks from samples of watermarked text raises the possibility of spoofing attacks. Successful execution of spoofing attacks suggests limitations for using watermarks to identify provenance, which has implications for the responsible usage and deployment of watermarking.
@@
Figure 1: A diagram illustrating decoding-based watermarking (top) vs. weights-based watermarking (bottom). Notice that decoding-based watermarking requires a specialized decoding algorithm $f_{\text{w}}$ to generate watermarked text, whereas weights-based watermarking generates watermarked text directly from the model, using just its weights. Watermark distillation enables weights-based watermarking by training a student model $p_{\theta}$ to behave like the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$.
@@
To enable weights-based watermarking, we propose logits-based and sampling-based watermark distillation, two simple methods for a student model to learn weights-based watermarking from a teacher model with decoding-based watermarking. Intuitively, in logits-based watermark distillation, the student model is trained to match the next token distribution outputted by the teacher model using decoding-based watermarking. In sampling-based watermark distillation, the teacher model with decoding-based watermarking is first used to generate watermarked samples. Then, the student model is fine-tuned on these watermarked samples.
@@
We experiment with three distinct decoding-based watermarking strategies: KGW (Kirchenbauer et al., 2023a), Aar (Aaronson, 2023), and KTH (Kuditipudi et al., 2023), and various values for their hyperparameters. We find that watermarks and hyperparameter settings vary in their degree of learnability. For example, in each watermarking strategy, higher-distortion hyperparameter settings are successfully learned by both forms of watermark distillation (median p-values less than 0.0001). Lower-distortion watermarks and hyperparameter settings are more challenging and less sample efficient to learn, but not unlearnable, as the p-values are still noticeably smaller than random chance.
We demonstrate applications of these findings towards developing watermarks for open-source models and the possibility of spoofing attacks. We find that weights-based watermarking is robust to changes in decoding algorithm and is a promising first step towards developing watermarking for open models. However, these watermarking capabilities are removed by further fine-tuning on normal text, indicating that improving robustness to fine-tuning is an important remaining challenge. For spoofing, we show that sampling-based watermark distillation is applicable to the spoofing setting, raising the possibility of spoofing attacks. We demonstrate a proof-of-concept attack on instruction-following chat models to substantiate this possibility, which suggests limitations on using watermarks to assign provenance and blame to specific models.
@@ 2 Background and notation: decoding-based watermarking
We study autoregressive language models $p_{\mathrm{LM}}:\mathcal{V}^{\ast}\to\Delta(\mathcal{V})$ that map from a prefix string $x$ to a next token distribution over the vocabulary $\mathcal{V}$.
Informally, a decoding-based watermarking strategy uses a watermark key to modify the model’s original next token distribution into a new distribution for generating watermarked text, which has a watermark signal embedded. The watermark detection algorithm looks for this signal using the same watermark key. Formally, we define a decoding-based watermarking strategy to be a function
@@
Given some text $x$ and watermark key $\xi$, $f_{\text{d}}$ outputs a p-value with respect to the null hypothesis that $x$ is independent of $f_{\text{w}}$ with key $\xi$. Internally, $f_{\text{d}}$ computes a test statistic that measures the strength of the signal embedded by the watermark, then computes a p-value using the distribution of the test statistic under the null hypothesis. If the p-value is below a certain significance level, the null hypothesis is rejected and the text is detected as watermarked. Slightly imprecisely, rejecting the null means the text is detected as model-generated. [^2] This is slightly imprecise because model-generated text is not the only text that can be watermarked. For example, someone could write watermarked text by manually following the watermarking algorithm. However, for most practical use cases, such as detecting academic dishonesty, this minor imprecision is not an issue because either way, the user is doing something suspicious and unusual.
@@
In this paper, we use three decoding-based watermarking strategies: Algorithm 2 in Kirchenbauer et al. (2023a), the Gumbel softmax scheme in Aaronson (2023), and the exponential minimum sampling scheme in Kuditipudi et al. (2023). Using the authors’ names and initials, we refer to these as KGW, Aar, and KTH, respectively.
2.1 KGW
@@
In the KGW watermarking strategy (Kirchenbauer et al., 2023a), when generating the next token, the vocabulary is pseudorandomly split into a “green list” and “red list” by hashing the previous token. The green list contains watermark hyperparameter $\gamma\in(0,1)$ proportion of the vocabulary. Hyperparameter $\delta>0$ is added to the original log-probabilities of the green list tokens, which are then softmaxed. This procedure makes green list tokens more likely in watermarked text than in non-watermarked text. At detection time, if the proportion of green list tokens in a text is much greater than $\gamma$, then the p-value is small.
Formally, the KGW decoding-based watermarking strategy is defined as
$f_{\text{w}}^{\text{KGW}}\left({\bm{p}},x,\xi;\gamma,\delta\right)=\mathrm{softmax}\bigl{(}\log({\bm{p}})+\delta\cdot f_{\text{hash}}^{\text{KGW}}\left(x_{\mathrm{len}(x)};\xi,\gamma,|\mathcal{V}|\right)\bigr{)}$ (3)
where $f_{\text{hash}}^{\text{KGW}}$ is parameterized by key $\xi$ and hashes the previous token $x_{\mathrm{len}(x)}$ and returns $g\in\{0,1\}^{|\mathcal{V}|}$, which contains $\gamma\cdot|\mathcal{V}|$ ones and $(1-\gamma)\cdot|\mathcal{V}|$ zeros, encoding the green list. The KGW watermark detection function is
$f_{\text{d}}^{\text{KGW}}\left(x,\xi;\gamma\right)=1-F_{B}\left(\sum\nolimits_{t=2}^{\mathrm{len}(x)}f_{\text{hash}}^{\text{KGW}}\left(x_{t-1};\xi,\gamma,|\mathcal{V}|\right)_{x_{t}}\right)$ (4)
@@
where $F_{B}$ is the cumulative distribution function (CDF) for binomial distributed random variable $B\sim\mathrm{Bin}(\mathrm{len}(x)-1,\gamma)$. $f_{\text{w}}^{\text{KGW}}$ distorts model outputs by upweighting green list tokens, increasing perplexity of generated texts computed by a large model (Kirchenbauer et al., 2023a). Increasing the bias hyperparameter $\delta$ increases detectability, i.e., smaller p-values, but also increases distortion.
2.2 Aar
The Aar watermarking strategy (Aaronson, 2023) uses a pseudorandom hash function $f_{\text{hash}}^{\text{Aar}}$ parametrized by key $\xi$ that hashes the previous $k$ tokens and returns $r\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$. $r$ assigns a score to each token in the vocabulary, and $k\in\mathbb{N}$ is a hyperparameter. The Aar decoding-based watermarking strategy is defined as
where $\mathrm{onehot}(i,|\mathcal{V}|)$ returns a $|\mathcal{V}|$-dimensional probability vector with $1$ at index $i$ and 00 everywhere else. Aar chooses the token that maximizes $r_{i}^{1/p_{i}}$, i.e., a token $i$ with both a high original probability $p_{i}$ and hash score $r_{i}$. The Aar watermark detection function is
@@ 2.2 Aar
where $F_{G}$ is the CDF for gamma distributed random variable $G\sim\mathrm{Gamma}(\mathrm{len}(x)-k,1)$. Because $f_{\text{w}}^{\text{Aar}}$ is deterministic based on the previous $k$ tokens and the original model probabilities, Aar can sometimes produce repetitive text, especially at small $k$ (Kuditipudi et al., 2023). Increasing $k$ decreases repetition as $k$-grams become less likely to be repeated, but the watermark also becomes less robust to edits, as each token edit affects the hashed watermark scores for $k+1$ tokens.
2.3 KTH
The KTH watermarking strategy (Kuditipudi et al., 2023) is similar to Aar, but instead of hashing the previous tokens to obtain the watermark scores $r$, the scores are obtained from the next position in the key sequence $\xi$. In KTH, $\xi=\xi^{(1)},\ldots,\xi^{(m)}$ where each $\xi^{(j)}\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$ and $m$ should be longer than the maximum generation length. Then, the KTH decoding-based watermarking strategy is defined as
@@ 2.3 KTH
where $\mathrm{onehot}$ is as defined for Aar. Note that $f_{\text{w}}^{\text{KTH}}$ is deterministic. To allow the model to generate different responses to the same prompt, before generating each sequence, $\xi$ can be shifted by some random $\tau$, i.e., $\xi^{\prime}=\xi^{(1+\tau\bmod m)},\ldots,\xi^{(m+\tau\bmod m)}$ and $\xi^{\prime}$ is used for $f_{\text{w}}^{\text{KTH}}$ (Kuditipudi et al., 2023). To study the impact of these shifts on learnability, we introduce an additional hyperparameter $s\in[1,m]$ which restricts the range of the shift (e.g., $s=1$ means the original $\xi$ is always used, and $s=m$ means that any shift of $\xi$ can be used). [^3] We space the $s$ shifts evenly across $[1,m]$, e.g., the set of possible shifts $\tau$ is $\{i\cdot\lfloor m/s\rfloor:0\leq i<s\}$. Increasing $s$ expands the range of possible generations the model can output.
At detection time, the basic test statistic is $\phi(x,\xi)=\sum_{t=1}^{\mathrm{len}(x)}-\log(1-\xi^{(t)}_{x_{t}})$. However, to be robust to edits, the detection statistic computes a minimum Levenshtein distance using the basic test statistic as a cost (see Definition 5 in Kuditipudi et al. (2023)) and compares it to the expected distribution under the null (see Algorithms 5 and 6 in Kuditipudi et al. (2023)). [^4] This expected reference distribution has finite number of samples $T$, and the p-values obtained by comparing to this distribution are lower bounded by $\frac{1}{T+1}$.
@@ 3 Methods
3.1 Problem statement
Given a teacher language model $p_{\mathrm{LM}}$, decoding-based watermarking strategy $f_{\text{w}}$, corresponding detection algorithm $f_{\text{d}}$, and key $\xi$, the goal is to learn a student model $p_{\theta}$ whose sampling distribution naturally generates watermarked text. If $p_{\theta}$ generates text that attains small detection p-value $f_{\text{d}}(y,\xi)$ with probability similar to that of $p_{\mathrm{LM}}$ with $f_{\text{w}}$, then we say that $p_{\theta}$ has learned a weights-based watermarking strategy, since $p_{\theta}$ has learned to generate watermarked text using just its weights. See Figure 1 for an illustration of decoding-based versus weights-based watermarking.
Next, we present two methods for learning a weights-based watermarking strategy: logits-based watermark distillation and sampling-based distillation, which fall under the broader category of knowledge distillation (Hinton et al., 2015; Kim & Rush, 2016).
3.2 Logits-based watermark distillation
Informally, in logits-based watermark distillation (WD-L), we train the student model to behave as if it had a decoding-based watermarking strategy applied. We train the student model $p_{\theta}$ to match the next token distribution outputted by the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ and key $\xi$, i.e., $f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x_{<t}\right),x_{<t},\xi\right)$. Formally, the training objective is to minimize the KL divergence loss
$\mathcal{L}_{\text{WD-L}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(x)}D_{\mathrm{KL}}\left(f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x_{<t}\right),x_{<t},\xi\right)\,\|\,p_{\theta}\left(\cdot\mid x_{<t}\right)\right)$ (8)
where $\mathcal{D}$ is the training dataset and teacher model $p_{\mathrm{LM}}$ is frozen. This approach requires that $p_{\mathrm{LM}}$ and $p_{\theta}$ have the same tokenizer and vocabulary so that the logits can be aligned between the two models. It is also helpful if both $p_{\mathrm{LM}}$ and $p_{\theta}$ have the same model architecture, as in this case we can initialize $p_{\theta}$ to $p_{\mathrm{LM}}$. One advantage of this approach is that the ground truth next tokens $x_{t}$ from dataset $\mathcal{D}$ is not used in the loss function, and thus $\mathcal{D}$ does not need to be watermarked text; standard datasets containing non-watermarked human-generated text can be used.

### v3

@@ Abstract
Watermarking of language model outputs enables statistical detection of model-generated text, which can mitigate harms and misuses of language models. Existing watermarking strategies operate by altering the decoder of an existing language model. In this paper, we ask whether language models can directly learn to generate watermarked text, which would have significant implications for the real-world deployment of watermarks. First, learned watermarks could be used to build open models that naturally generate watermarked text, enabling watermarking for open models, where users can control the decoding procedure. Second, if watermarking is used to determine the provenance of generated text, an adversary can hurt the reputation of a victim model by spoofing its watermark and generating damaging watermarked text. To investigate the learnability of watermarks, we propose watermark distillation, which trains a student model to behave like a teacher model that uses decoding-based watermarking. We test our approach on three decoding-based watermarking strategies and various hyperparameter settings, finding that models can learn to generate watermarked text with high detectability. We also find limitations to learnability, including the loss of watermarking capabilities under fine-tuning on normal text and high sample complexity when learning low-distortion watermarks. [^1] See https://github.com/chenchenygu/watermark-learnability for code and models.
@@ 1 Introduction
As language models (LMs) become more capable and widely used, watermarking LM outputs becomes increasingly important to mitigate potential harms and misuses of LMs. Watermarking enables statistical detection of LM-generated text, which enables enforcing policies on LM usage, e.g., removing LM-generated disinformation from social media platforms or detecting academic dishonesty. Another proposed use case of watermarking is identifying the provenance of text, i.e., tracing text to the specific LM that generated it (Abdelnabi & Fritz, 2021; Kuditipudi et al., 2023).
@@
Recent works have shown that it is possible for an LM provider to inject specific, known watermark signals into text using specialized decoding algorithms (Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023), but little is known about whether these watermarks are learnable by a model. The learnability of watermarks has significant implications for the real-world deployment of watermarks, as it could enable downstream applications and adversarial spoofing attacks.
@@
In this work, we study the learnability of watermarks by studying weights-based watermarking, which involves learning parameters for a language model that cause it to generate watermarked text under its natural sampling distribution, without using a special decoding-time watermarking algorithm. Our investigation is motivated by its relevant implications for two applications: (i) developing watermarking for open language models and (ii) spoofing watermarks.
First, existing watermarking methods depend upon using a specialized decoding algorithm, making them too inflexible for open LMs. For open LMs, where the weights are released, a user can use an ordinary decoding algorithm and generate non-watermarked text, whether intentionally or not. We find that weights-based watermarking works with standard decoding strategies, removing the reliance on a specialized decoder. This makes it a promising first step towards developing watermarking for open LMs. However, we also find that weights-based watermarking capabilities can be removed by fine-tuning on normal text, indicating that improving robustness to fine-tuning is an important remaining challenge.
Second, in watermark spoofing attacks, an adversary outputs text that contains the watermark signal from a victim LM (Sadasivan et al., 2023). If watermarking is used to identify the provenance of text, then an attacker could attribute damaging text to the victim LM and hurt its reputation. We find that the learning of weights-based watermarking can enable spoofing attacks, and we demonstrate a proof-of-concept attack on an instruction-following chat model. The possibility of spoofing attacks suggests that watermarking should not be used to attribute provenance or blame to a specific LM. Instead, watermarking should only be used to statistically detect LM-generated text, which can be used for tasks such as finding infractions of policies on LM usage.
@@
Figure 1: Decoding-based watermarking (top) versus weights-based watermarking (bottom). Decoding-based watermarking requires a specialized decoding algorithm $f_{\text{w}}$ to generate watermarked text, whereas weights-based watermarking can use standard decoding to generate watermarked text directly from the model, using just its weights. Watermark distillation enables weights-based watermarking by training a student model $p_{\theta}$ to behave like the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$.
@@
To enable weights-based watermarking, we propose logit-based and sampling-based watermark distillation, two simple methods for a student model to learn weights-based watermarking from a teacher model with decoding-based watermarking. Intuitively, in logit-based watermark distillation, the student model is trained to match the next token distribution outputted by the teacher model using decoding-based watermarking. In sampling-based watermark distillation, the teacher model with decoding-based watermarking is first used to generate watermarked samples. Then, the student model is fine-tuned on these watermarked samples.
@@
We experiment with three decoding-based watermarking strategies: KGW (Kirchenbauer et al., 2023a; Zhao et al., 2023a), Aar (Aaronson, 2023), and KTH (Kuditipudi et al., 2023), and various values for their hyperparameters that control the level of distortion induced by watermarking. We find that watermarks and hyperparameter settings vary in their degree of learnability. In each watermarking strategy, higher-distortion hyperparameter settings are successfully learned by both forms of watermark distillation (median p-values less than 0.0001). Lower-distortion watermarks and hyperparameter settings are more challenging and less sample efficient to learn, but not unlearnable, as the p-values are still noticeably smaller than the non-watermarked baseline of 0.5.
@@ 2 Background and notation: decoding-based watermarking
We study autoregressive language models $p_{\mathrm{LM}}:\mathcal{V}^{\ast}\to\Delta(\mathcal{V})$ that map from a prefix string $x\in\mathcal{V}^{\ast}$ to a next token distribution over the vocabulary $\mathcal{V}$. Informally, a decoding-based watermarking strategy $f_{\text{w}}$ uses a watermark key $\xi$ to modify the model’s original next token distribution $p_{\mathrm{LM}}(\cdot\mid x)$ into a new distribution for generating watermarked text, which has a watermark signal embedded. The watermark detection algorithm $f_{\text{d}}$ looks for this signal using the same watermark key $\xi$.
Formally, we define a decoding-based watermarking strategy to be a function
@@
Given some text $x$ and watermark key $\xi$, $f_{\text{d}}$ outputs a p-value with respect to the null hypothesis that $x$ is independent of $f_{\text{w}}$ with key $\xi$. Informally, $f_{\text{d}}$ computes a test statistic that measures the strength of the watermark signal, then computes a p-value using the distribution of the test statistic under the null hypothesis. If the p-value is below a given significance level, the null hypothesis is rejected and the text is detected as watermarked. Slightly imprecisely, rejecting the null hypothesis means the text is detected as model-generated. [^2] This is slightly imprecise because model-generated text is not the only text that can be deliberately watermarked. For example, a human could write watermarked text by manually following the watermarking algorithm. However, for most practical use cases, such as detecting academic dishonesty, this minor imprecision is not an issue because either way, the user is doing something suspicious and unusual.
@@
In this paper, we consider three decoding-based watermarking strategies: Algorithm 2 in Kirchenbauer et al. (2023a), the Gumbel softmax scheme in Aaronson (2023), and the exponential minimum sampling scheme in Kuditipudi et al. (2023). Using the authors’ names and initials, we refer to these as KGW, Aar, and KTH, respectively. We briefly describe these watermarking strategies below. See Appendix A for additional details and formal definitions.
2.1 KGW: green list bias
@@
In the KGW watermarking strategy (Kirchenbauer et al., 2023a), when generating the next token, the vocabulary is pseudorandomly split into a “green list” and “red list” by hashing the previous token using the watermark key $\xi$. The green list contains watermark hyperparameter $\gamma\in(0,1)$ proportion of the vocabulary. Then, before the model’s logits are converted to probabilities via the softmax function, hyperparameter $\delta>0$ is added to the logits of the green list tokens. This procedure makes green list tokens more likely in watermarked text than in non-watermarked text. So, at detection time, if the proportion of green list tokens in a text is much greater than $\gamma$, then the p-value is small.
More generally, the previous $k$ tokens can be hashed, where $k$ is a hyperparameter. Values of $k>1$ are investigated by Kirchenbauer et al. (2023b), finding that lower $k$ leads to more repetitive outputs. When $k=0$, the green and red lists are fixed, regardless of the previous tokens. $k=0$ was proposed by Zhao et al. (2023a) as Unigram-Watermark, a variant of KGW, but we will denote it as KGW $k=0$ to simplify notation.
@@
KGW distorts model outputs by upweighting green list tokens, increasing perplexity of generated texts computed by a larger model (Kirchenbauer et al., 2023a). Increasing the bias hyperparameter $\delta$ increases detectability, i.e., smaller p-values, but also increases distortion.
2.2 Aar: boosting continuous hash scores
The Aar watermarking strategy (Aaronson, 2023) hashes the previous $k$ tokens using key $\xi$ (where $k$ is a hyperparameter) to obtain a score $r_{i}$ for each token $i\in\mathcal{V}$, where each $r_{i}$ is uniformly distributed in $[0,1]$. Let $p_{i}$ be the original model probability for token $i$. Then, the next generated token is deterministically chosen to be the token $i$ which maximizes $r_{i}^{1/p_{i}}$, i.e., a token with both a high original probability $p_{i}$ and high hash score $r_{i}$. This procedure boosts the hash scores of tokens in watermarked text compared to non-watermarked text. So, at detection time, if the hash scores $r_{i}$ of the tokens in the observed sequence are high, then the p-value is low.
@@ 2.2 Aar: boosting continuous hash scores
Since Aar deterministically selects the next token based on the previous $k$ tokens and the original model probabilities, Aar can lead to repetitive text, especially for small $k$ (Kuditipudi et al., 2023). Increasing $k$ decreases repetitiveness, as larger $k$-grams are less likely to be repeated, but the watermark also becomes less robust to edits, as each token edit affects the hash scores for $k+1$ tokens.
2.3 KTH: robust sequence alignment
@@
The KTH watermarking strategy (Kuditipudi et al., 2023) is similar to Aar, but instead of hashing previous tokens to obtain the scores $r_{i}$, the scores are obtained from the next element in the key sequence $\xi$. In KTH, $\xi=\xi^{(1)},\ldots,\xi^{(m)}$ where each $\xi^{(j)}\in[0,1]^{|\mathcal{V}|}$ contains the scores, with entries uniformly distributed across $[0,1]$. Then, to generate the $j$-th token in the sequence, KTH deterministically chooses the token $i$ that maximizes $\left(\xi^{(j)}_{i}\right)^{1/p_{i}}$. Note that $m$ should be larger than the maximum generation length. To allow different generations from the same prompt, before generating each sequence, $\xi$ can be shifted by some random $\tau$, i.e., $\xi^{\prime}=\xi^{(1+\tau\bmod m)},\ldots,\xi^{(m+\tau\bmod m)}$. To study the impact of these shifts on learnability, we introduce a hyperparameter $s\in[1,m]$ for how many shift values $\tau$ are possible. [^3] We space the $s$ shifts evenly across $[1,m]$, i.e., the set of possible shifts $\tau$ is $\{i\cdot\lfloor m/s\rfloor:0\leq i<s\}$. Increasing $s$ expands the range of possible model generations.
At detection time, to be robust to text edits and shifts, the test statistic quantifies how well a text $x$ can be aligned with the key sequence $\xi$. More specifically, the test statistic computes a minimum Levenshtein distance using the alignment cost $d(x,\xi)=\sum_{t=1}^{\mathrm{len}(x)}\log(1-\xi^{(t)}_{x_{t}})$. A lower (more negative) test statistic indicates stronger watermark signal. To compute p-values, the observed test statistic is compared to a reference distribution of test statistics of non-watermarked texts. Letting $T$ be the number of samples in the reference distribution, the p-values computing using this method are lower bounded by $\frac{1}{T+1}$.
@@ 3 Methods
Problem statement. Given a teacher model $p_{\mathrm{LM}}$, decoding-based watermarking strategy $f_{\text{w}}$, and key $\xi$, the goal is to learn a student model $p_{\theta}$ whose sampling distribution naturally generates watermarked text. Specifically, letting $f_{\text{d}}$ be the detection algorithm corresponding to $f_{\text{w}}$, if $p_{\theta}$ generates text $y$ with small detection p-value $f_{\text{d}}(y,\xi)$ with probability similar to that of $p_{\mathrm{LM}}$ with $f_{\text{w}}$, then $p_{\theta}$ has learned a weights-based watermarking strategy, since $p_{\theta}$ has learned to generate watermarked text using just its weights. Figure 1 illustrates decoding-based versus weights-based watermarking.
Next, we present two methods for learning a weights-based watermarking strategy: logit-based watermark distillation and sampling-based watermark distillation, which fall under the broader category of knowledge distillation (Hinton et al., 2015; Kim & Rush, 2016).
3.1 Logit-based watermark distillation
In logit-based watermark distillation, we train the student model $p_{\theta}$ to behave as if it had decoding-based watermarking strategy $f_{\text{w}}$ applied. Specifically, given an input $x$, we want the student model’s next token distribution $p_{\theta}(\cdot\mid x)$ to match $f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x\right),x,\xi\right)$, the next token distribution outputted by the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ and key $\xi$. So, given a dataset of texts $\mathcal{D}$, the training objective is to minimize the mean KL divergence between the teacher and student next token distributions on all prefixes in $\mathcal{D}$, given by
$\mathcal{L}_{\mathrm{logit}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(x)}D_{\mathrm{KL}}\left(f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x_{<t}\right),x_{<t},\xi\right)\,\|\,p_{\theta}\left(\cdot\mid x_{<t}\right)\right).$ (3)
The teacher model $p_{\mathrm{LM}}$ is frozen. This approach requires that $p_{\mathrm{LM}}$ and $p_{\theta}$ have the same tokenizer and vocabulary so that the logits can be aligned between the two models. It is also helpful if $p_{\mathrm{LM}}$ and $p_{\theta}$ share the same model architecture, as then we can initialize $p_{\theta}$ to $p_{\mathrm{LM}}$. Note that the ground truth next tokens $x_{t}$ from dataset $\mathcal{D}$ are not used in the loss function, so $\mathcal{D}$ does not need to be watermarked text. Standard datasets containing non-watermarked human-generated text can be used. [^4] If $\mathcal{D}$ is non-watermarked text, then it theoretically might be out of distribution for $p_{\theta}$ to autoregressively generate watermarked text, since $p_{\theta}$ would be conditioning on the watermarked text it has already generated. However, empirically, we find that logit-based distilled models can learn to generate watermarked text.